the area A of a rectangle increased by 4 is equal to 36 decreased by 5 times the area. what is the area of the rectangle? with equation and solution
Answer by cleomenius(959)

x+ 4 = 36 - 5x.
6x = 32
x = 5.33
5.33 + 4 = 9.33
36 - 5(5.33) = 9.35 (round off error)
Cleomenius.